OpenStudy (anonymous):

Answer the following probability questions: 1. There are 52 cards in a deck. There are 4 tens. What is the probability that if you cut the deck, you'll get a ten? Express your answer as a fraction reduced to its simplest form. 2. There are 52 cards in a deck. There are 4 eights and 4 nines. What is the probability that if you cut the deck, you'll get an eight or a nine? Express your answer as a decimal number rounded to the hundredths place. 3. There are 52 cards in a deck. There are 13 different cards in the deck that are clubs. What is the probability that if you cut the deck, you'll

OpenStudy (dumbcow):

probability is desired outcomes over total outcomes 1. 4/52 = 1/13 2. 8/52 = 2/13 3. 13/52 = 1/4

OpenStudy (henryblah):

1) 4 out of 52 cards are a 10. Therefore 1/13 2) 8 out of 52. Therefore 2/13. Use a calc to find decimal 3) 13 out of 52. Therefore 1/4

OpenStudy (anonymous):

Okay. Thanks. But what are 4 &5?

OpenStudy (henryblah):

Can't see 4 and 5...